# LLM-based Test-driven Interactive Code Generation: User Study and Empirical Evaluation

### v1

@@ Abstract
In this paper, we propose a novel interactive workflow TiCoder for guided intent clarification (i.e., partial formalization) through tests to support the generation of more accurate code suggestions. Through a mixed methods user study with 15 programmers, we present an empirical evaluation of the effectiveness of the workflow to improve code generation accuracy. We find that participants using the proposed workflow are significantly more likely to correctly evaluate AI generated code, and report significantly less task-induced cognitive load. Furthermore, we test the potential of the workflow at scale with four different state-of-the-art LLMs on two python datasets, using an idealized proxy for a user feedback. We observe an average absolute improvement of 38.43% in the pass@1 code generation accuracy for both datasets and across all LLMs within 5 user interactions, in addition to the automatic generation of accompanying unit tests.
@@ I Introduction
Large Language Models (LLMs) have shown tremendous potential in generating natural-looking programs from informal intent expressed in natural language. There has been a surge in research around training LLMs over programming language artifacts in just the last couple of years [Chen et al.(2021), Chowdhery et al.(2022), Nijkamp et al.(2022), Fried et al.(2022), Xu et al.(2022a)]. Commercial offerings such as GitHub Copilot [GitHub(2022)] are widely available, and have been shown to generate a non-trivial fraction of code in real-world scenarios [Ziegler et al.(2022)].
However, there are several challenges that arise when generating code from natural language specifications.[Liang et al.(2023), Xu et al.(2022b)]. For example, natural language prompts crafted by users may not always fully capture a their intent, as they may contain ambiguous language and lack of nuance. More importantly, it is not possible to automatically evaluate whether code generated from a natural language prompt is correct. Natural language is inherently ambiguous and enforcing the user intent through some mechanical process (such as testing, static analysis or formal verification) is not immediately possible.
Consider the following docstring, taken from MBPP [Austin et al.(2021b)], a popular Python programming tasks benchmark:
@@
While the intent may seem obvious at first, it is not immediately clear how to check the correctness of a potential solution. Querying an LLM such as text-davinci-003 [Ouyang et al.(2022)] yields several plausibly correct code implementations that pass simple tests such as rejecting the empty string “ ”, or accepting the string "aa_bb". However, it may also produce subtly buggy code solutions that accept strings such as "aa_bb_cc", which is inconsistent with the original user intent that expects the string to consist entirely of two sequences of lowercase letters joined by an underscore (as defined by the accompanying hidden reference solution and the validation tests from MBPP). In practice, this can often lead to users accepting code with subtle bugs while using LLMs [Asare et al.(2022), Perry et al.(2022)]. The apparent ambiguity in this particular docstring, and more importantly the informal nature of natural language, highlights the inability to immediately ascertain the correctness of the code generated by an LLM. Instead, it would be desirable to avoid surfacing such subtly incorrect codes by first clarifying, and partially formalizing, the user intent into a checkable specification.
This issue can be compounded when users are presented with a list of candidate suggestions from LLMs, such as in the Copilot VSCode IDE suggestions pane, which can display up to 10 suggestions. Users often have to linearly scan the list of code suggestions, review them, and reject the incorrect ones until arriving at one that satisfies their intent. In such situations, subtle bugs may be overlooked, with significant downstream impacts. In fact, several recent works exploring developer-AI interaction have highlighted the need for mechanisms to facilitate verification of AI-generated code[Bird et al.(2022), Xu et al.(2022b)], such as those that allow users to use tests that disambiguate between the different code suggestions [Jha et al.(2010)]. However, prior research has shown that it can be difficult for users to manually provide a sufficient number of test cases to disambiguate suggestions upfront [Lau(2009)].
Inspired by findings around example generation and disambiguation techniques in Programming By Examples (PBE) [Zhang et al.(2020)], and recent emerging ability of LLMs to generate tests [Lemieux et al.(2023), Dinella et al.(2022), Schäfer et al.(2023)] in this paper, we propose leveraging user-feedback through LLM-generated tests to improve correctness of LLM-generated code. Specifically, we propose the workflow of test-driven interactive code generation (TiCoder) to (a) clarify (i.e., partially formalize) user intent through generated tests, and (b) generate a ranked list of code that is consistent with such tests.
@@
In the case of LLM-based code generation, the generated tests not only help make natural language intent more precise and prune incorrect suggestions generated by the LLM, but can also serves as debugging aid for remaining suggestions and regression tests for future code edits [Ziegler et al.(2022)].
While the proposed framework appears intuitive, the utility of the interactive framework is contingent upon (a) the ability of LLMs to generate useful tests, and (b) the cost-benefit trade-off of the overhead of user interaction versus the benefit on pruning and ranking of code suggestions. To this end, we seek to understand: How does the proposed workflow impact the performance of developers evaluating AI generated code? In addition, the proposed framework should scale, augmenting the code generation accuracy of several open and closed-source LLMs. Thus we also seek to answer: Does proposed workflow augment the accuracy of code generation models?
@@
    We further evaluate the performance of the TiCoder workflow at scale by simulating user feedback, using the reference code solution as an idealized proxy. TiCoder is evaluated on on two Python datasets, MBPP and HumanEval, and a mixture of four open and closed sourced LLMs. We demonstrate that TiCoder contributes to improving the code generation accuracy of all LLMs considered. We observe an average absolute improvement of 38.43% in pass@1 code generation accuracy within 5 user interactions across both benchmarks. For MBPP, we observe an increase in pass@1 metric by an average of $23.88\%$ with a single user query, and by $41.60\%$ with 5 user queries. For HumanEval pass@1 increases by an average of $15.25\%$ with a single user query, and by $35.26\%$ with 5 user queries.
@@ II-1 Improving Code Generation Accuracy
Techniques for improving code generation accuracy is a rapidly growing field of work. Unlike the work proposed in this paper, these techniques do not consider user feedback, or guide users in clarifying their intent formally; we cover them briefly. AlphaCode [Li et al.(2022)] and CodeT [Chen et al.(2022)] both propose techniques to improve code generation accuracy by generating tests using LLMs, and then grouping code suggestions by the set of tests that they satisfy. CodeT [Chen et al.(2022)] refines the approach by scoring tests and code suggestions simultaneously by prioritizing tests that satisfy many code suggestions and prioritizing codes that satisfy many tests. As part of future work, we plan to explore if our approach may benefit from code and test ranking algorithms in CodeT. Similarly, work on program synthesis [Gulwani et al.(2017), Solar-Lezama(2009)] generates code that satisfies a formal specification either expressed as a logical specification or input-output tests [Gulwani(2011)]. Our work differs in that we use LLMs to generate code from informal specifications, i.e. natural language intent. However, it would be interesting for future work to leverage user-provided tests to improve the quality of code generation, as explored in recent works [Jain et al.(2022), Rahmani et al.(2021)]. In this work, to evaluate our proposed approach at scale, we simulate user feedback using the code reference implementation as an idealized proxy, similar to prior works in oracle-guided inductive synthesis [Jha et al.(2010), Jha and Seshia(2017)] and interactive program synthesis [Le et al.(2017), Ji et al.(2020)] where an an oracle (reference implementation or users) is queried to identify the output for a given input. However, prior works in this area appeal to an automatic symbolic engine (such as a constraint solver [Ji et al.(2020)] or automata construction [Zhang et al.(2020)]) to generate distinguishing example inputs for a pair of programs, which is inconceivable for general purpose imperative programming languages such as Python.
@@ II-2 Usability of AI Programming Assistants
Liang et al. [Liang et al.(2023)] identify that giving up on incorporating generated code, and lack of ability to provide feedback, are the most common usability issues encountered when using completion-based AI programming assistants. This often occurs because the code does not implement the desired functionality, participants do not know why certain code was generated and had trouble controlling the output to be aligned with their desired intent.
McNutt et al. [McNutt et al.(2023)] enumerate a design space of interactions with code assistants, including how users should be able to disambiguate candidate programs or refine their initial specifications, echoing prior studies have indicated that disambiguation can be valuable in the context of assistants like GitHub Copilot [Barke et al.(2023)] and traditional program synthesis tools [Mayer et al.(2015)]. Similarly, Xu et al. [Xu et al.(2022b)] explored challenges of IDE-based AI assistants, including how well specified the queries that users formulate are. They find that participants frequently have trouble expressing intent in their natural language queries to the assistant, and issues of under specification often relate to ambiguous instructions, such as omitting variable names.
Mozannar et al. [Mozannar et al.(2022)] identify 12 core activities associated with using GitHub Copilot and find that programmers often iterate on their prompts until they obtain the suggestion they desire, and spend a significant amount of time verifying code suggestions. In fact, recent work by Bird et al. [Bird et al.(2022)] shows that as result of AI-powered tools, developer roles are shifting so that more time is spent time reviewing code than actually writing code. Our work builds upon observations of previous studies, and explores mechanisms to support code evaluation tasks.
@@ IV-A High-level Workflow
We present both the scenarios as they enjoy complementary benefits. The TiCoder-PassFail scenario is more lightweight, in terms of user feedback, as well as, generalizes well for richer tests beyond input-output examples. For example, tests for stateful APIs comprises of a test-prefix as input and the output oracle consists of a non-trivial predicate (e.g., checking functional correctness of a stack object using the predicate s.pop() == a on a stack object s and element a[Dinella et al.(2022)]. On the other hand, TiCoder-Output puts less burden on an LLM to create the correct output for a given test input; relying instead on the user. However, it may require the user to specify a possibly non-trivial test oracle when used beyond input-output examples.
@@ IV-B2 Ranking test suggestions
After obtaining the set of tests produced by an LLM, the user is presented with a sequence of tests. The user response to these proposed tests in both TiCoder scenarios (TiCoder-PassFail, TiCoder-Output) are used to prune and rank code suggestions. To minimize the number of user interactions, it is desirable to prioritize tests that would result in the most number of incorrect code suggestions being pruned away [Jha et al.(2010), Le et al.(2017)]. To achieve this, the set of tests are executed against the set of possible code suggestions generated by the LLM.
Then, using this execution information, we adopt a discriminative test ranking policy that prioritizes tests that can discriminate best among the set of code suggestions generated by the LLM.
If a test $t$ can discriminate between code suggestions well (i.e., splits the set of code suggestions into roughly equal halves), then it would prune away a substantial fraction of the code suggestions irrespective of the user response (either Pass or Fail).
More precisely, let $U$ be the set of test suggestions and $G$ be the set of code suggestions that have not been pruned away after $k\geq 0$ user interactions. For each test $t\in U$, we split the set of code suggestions $G$ into the sets $G^{+}_{t}$ and $G^{-}_{t}$ of code suggestions that pass and fail the assertion in $t$, respectively. We then prioritize tests where the ratio of the sizes of these two set is closest to 1. In other words, we rank the tests in decreasing order using the following scoring metric $s_{discr}$:
@@
Consider the example in Figure 3. Consider the two tests $t_{1}$ and $t_{2}$: Two code suggestions {$c_{2}$, $c_{3}$} FAIL on test suggestion $t_{1}$ while one suggestion {$c_{1}$} PASS, making $s_{\textit{discr}}(t_{1})=\texttt{min}(1,2)/\texttt{max}(1,2)=1/2$. Similarly, two code suggestions {$c_{1}$, $c_{2}$} PASS on test suggestion $t_{2}$ while one suggestion {$c_{3}$} FAIL and $s_{\textit{discr}}(t_{2})=1/2$.
All code suggestions in this example PASS on test $t_{3}$ making $s_{\textit{discr}}(t_{3})=0$.
@@ IV-B3 Pruning and ranking code suggestions
TiCoder returns a ranked list of code suggestions, whose behavior is consistent with all the user responses, and prunes the other code suggestions generated by the LLM, whose execution behaviour on tests is contradictory to user expectation. Let us first consider the case of code pruning. Let $t\doteq(i,o)$ be a test in the form of an input-output example presented to the user. If the user responds Pass, then we prune any code $c\in G$ for which executing $c(i)\neq o$. Similarly, if the user responds Fail, then we prune any code $c\in G$ for which executing $c(i)=o$. In addition, for TiCoder-Output if the user provides the desired output $o^{\prime}$ for the input $i$, then we can further prune any code suggestion $c$ for which $c(i)\neq o^{\prime}$. Note that we cannot soundly prune any code if the user responds with Undefined.
@@ V-A1 Control condition: AI Programming Assistant 1
Assistant 1 represents the control condition for the experiment. Given the pre-selected prompt, Assistant 1 generates 5 code suggestions for the user, surfaced in a random order. Participants using Assistant 1 always see 5 unique code suggestions. We make this decision to reflect the current user experience scenario of several real-world AI code generation tools, such as GitHub Copilot’s completion panel. For example, the GitHub Copilot completion panel in VSCode shows the user up to 10 possible code suggestions at a time. Our decision is also informed by research pointing to the benefit of surfacing multiple code suggestions[McNutt et al.(2023), Mayer et al.(2015), Barke et al.(2023), Liang et al.(2023)]. We limit the maximum number of codes to 5 so as to allow the participant to complete each task within 15 minutes.

### v2

@@ Abstract
In this paper, we propose a novel interactive workflow TiCoder for guided intent clarification (i.e., partial formalization) through tests to support the generation of more accurate code suggestions. Through a mixed methods user study with 15 programmers, we present an empirical evaluation of the effectiveness of the workflow to improve code generation accuracy. We find that participants using the proposed workflow are significantly more likely to correctly evaluate AI generated code, and report significantly less task-induced cognitive load. Furthermore, we test the potential of the workflow at scale with four different state-of-the-art LLMs on two python datasets, using an idealized proxy for a user feedback. We observe an average absolute improvement of 45.97% in the pass@1 code generation accuracy for both datasets and across all LLMs within 5 user interactions, in addition to the automatic generation of accompanying unit tests.
@@ I Introduction
Large Language Models (LLMs) have shown tremendous potential in generating natural-looking programs from informal intent expressed in natural language. There has been a surge in research around training LLMs over programming language artifacts in just the last couple of years [1, 2, 3, 4, 5]. Commercial offerings such as GitHub Copilot [6] are widely available, and have been shown to generate a non-trivial fraction of code in real-world scenarios [7].
However, there are several challenges that arise when generating code from natural language specifications.[8, 9]. For example, natural language prompts crafted by users may not always fully capture a their intent, as they may contain ambiguous language and lack of nuance. More importantly, it is not possible to automatically evaluate whether code generated from a natural language prompt is correct. Natural language is inherently ambiguous and enforcing the user intent through some mechanical process (such as testing, static analysis or formal verification) is not immediately possible.
Consider the following docstring, taken from MBPP [10], a popular Python programming tasks benchmark:
@@
While the intent may seem obvious at first, it is not immediately clear how to check the correctness of a potential solution. Querying an LLM such as text-davinci-003 [11] yields several plausibly correct code implementations that pass simple tests such as rejecting the empty string “ ”, or accepting the string "aa_bb". However, it may also produce subtly buggy code solutions that accept strings such as "aa_bb_cc", which is inconsistent with the original user intent that expects the string to consist entirely of two sequences of lowercase letters joined by an underscore (as defined by the accompanying hidden reference solution and the validation tests from MBPP). In practice, this can often lead to users accepting code with subtle bugs while using LLMs [12, 13]. The apparent ambiguity in this particular docstring, and more importantly the informal nature of natural language, highlights the inability to immediately ascertain the correctness of the code generated by an LLM. Instead, it would be desirable to avoid surfacing such subtly incorrect codes by first clarifying, and partially formalizing, the user intent into a checkable specification.
This issue can be compounded when users are presented with a list of candidate suggestions from LLMs, such as in the Copilot VSCode IDE suggestions pane, which can display up to 10 suggestions. Users often have to linearly scan the list of code suggestions, review them, and reject the incorrect ones until arriving at one that satisfies their intent. In such situations, subtle bugs may be overlooked, with significant downstream impacts. In fact, several recent works exploring developer-AI interaction have highlighted the need for mechanisms to facilitate verification of AI-generated code[14, 9], such as those that allow users to use tests that disambiguate between the different code suggestions [15].
However, prior research has shown that it can be difficult for users to manually provide a sufficient number of test cases to disambiguate suggestions upfront [16].
Inspired by findings around example generation and disambiguation techniques in Programming By Examples (PBE) [17], and recent emerging ability of LLMs to generate tests [18, 19, 20] in this paper, we propose leveraging user-feedback through LLM-generated tests to improve the trust and correctness of LLM-generated code. Specifically, we propose the workflow of test-driven interactive code generation (TiCoder) to (a) clarify (i.e., partially formalize) user intent through generated tests, and (b) generate a ranked list of code that is consistent with such tests.
@@
In the case of LLM-based code generation, the generated tests not only help make natural language intent more precise and prune incorrect suggestions generated by the LLM, but can also serves as debugging aid for remaining suggestions and regression tests for future code edits [7].
While the proposed framework appears intuitive, it may not scale to more complex code generation tasks. For example, in cases where the user is unable to validate tests, e.g. for tests that require intricate testing frameworks, the a workflow may not be tenable. Furthermore, the utility of the interactive framework is contingent upon (a) the ability of LLMs to generate useful tests, and (b) the cost-benefit trade-off of the overhead of user interaction versus the benefit on pruning and ranking of code suggestions.
To this end, we seek to understand: How does the proposed workflow impact the performance of developers evaluating AI generated code? In addition, the proposed framework should scale, augmenting the code generation accuracy of several open and closed-source LLMs. Thus we also seek to answer: Does proposed workflow augment the accuracy of code generation models?
@@
    We further evaluate the performance of the TiCoder workflow at scale by simulating user feedback, using the reference code solution as an idealized proxy. TiCoder is evaluated on on two Python datasets, MBPP and HumanEval, and a mixture of four open and closed sourced LLMs. We demonstrate that TiCoder contributes to improving the code generation accuracy of all LLMs considered. We observe an average absolute improvement of 45.73% in pass@1 code generation accuracy within 5 user interactions across both benchmarks. In fact, we observe TiCoder can boost smaller model pass@1 accuracy to levels comparable to much larger models, such as GPT-4-32k, within just one user interaction.
@@ II-1 Improving Code Generation Accuracy
Techniques for improving code generation accuracy is a rapidly growing field of work. Unlike the work proposed in this paper, these techniques do not consider user feedback, or guide users in clarifying their intent formally; we cover them briefly.
AlphaCode [21] and CodeT [22] both propose techniques to improve code generation accuracy by generating tests using LLMs, and then grouping code suggestions by the set of tests that they satisfy. CodeT [22] refines the approach by scoring tests and code suggestions simultaneously by prioritizing tests that satisfy many code suggestions and prioritizing codes that satisfy many tests. While there are similarities with CodeT in using LLM generated tests to rerank generated code that results in code generation accuracy on benchmarks, TiCoder is complementary as one can apply TiCoder after CodeT. But more importantly, we argue that a user cannot trust the generated code from CodeT any more than using LLM directly. This is because the user is still presented with a set of code suggestions. In contrast, with TiCoder, we first formalize the user intent through tests allowing the user to constrain the code that the user will need to eventually sample from. TiCoder also allows users to modify test output in one setting, which is not possible in the CodeT approach, where the tests are fixed throughout. As part of future work, we plan to explore if our approach may benefit from code and test ranking algorithms in CodeT.
Similarly, work on program synthesis [23, 24] generates code that satisfies a formal specification either expressed as a logical specification or input-output tests [25]. Our work differs in that we use LLMs to generate code from informal specifications, i.e. natural language intent. However, it would be interesting for future work to leverage user-provided tests to improve the quality of code generation, as explored in recent works [26, 27]. In this work, to evaluate our proposed approach at scale, we simulate user feedback using the code reference implementation as an idealized proxy, similar to prior works in oracle-guided inductive synthesis [15, 28] and interactive program synthesis [29, 30] where an an oracle (reference implementation or users) is queried to identify the output for a given input. However, prior works in this area appeal to an automatic symbolic engine (such as a constraint solver [30] or automata construction [17]) to generate distinguishing example inputs for a pair of programs, which is inconceivable for general purpose imperative programming languages such as Python.
@@ II-2 Usability of AI Programming Assistants
Liang et al. [8] identify that giving up on incorporating generated code, and lack of ability to provide feedback, are the most common usability issues encountered when using completion-based AI programming assistants. This often occurs because the code does not implement the desired functionality, participants do not know why certain code was generated and had trouble controlling the output to be aligned with their desired intent.
McNutt et al. [31] enumerate a design space of interactions with code assistants, including how users should be able to disambiguate candidate programs or refine their initial specifications, echoing prior studies have indicated that disambiguation can be valuable in the context of assistants like GitHub Copilot [32] and traditional program synthesis tools [33]. Similarly, Xu et al. [9] explored challenges of IDE-based AI assistants, including how well specified the queries that users formulate are. They find that participants frequently have trouble expressing intent in their natural language queries to the assistant, and issues of under specification often relate to ambiguous instructions, such as omitting variable names.
Mozannar et al. [34] identify 12 core activities associated with using GitHub Copilot and find that programmers often iterate on their prompts until they obtain the suggestion they desire, and spend a significant amount of time verifying code suggestions. In fact, recent work by Bird et al. [14] shows that as result of AI-powered tools, developer roles are shifting so that more time is spent time reviewing code than actually writing code. Several recent works[35, 36, 37] identify clear opportunities for improving the accuracy of LLM code generation techniques. Our work builds upon observations of previous studies, and explores mechanisms to support code evaluation tasks.
@@ IV-A High-level Workflow
We present both the scenarios as they enjoy complementary benefits. The TiCoder-PassFail scenario is more lightweight, in terms of user feedback, as well as, generalizes well for richer tests beyond input-output examples. For example, tests for stateful APIs comprises of a test-prefix as input and the output oracle consists of a non-trivial predicate (e.g., checking functional correctness of a stack object using the predicate s.pop() == a on a stack object s and element a[19]. On the other hand, TiCoder-Output puts less burden on an LLM to create the correct output for a given test input; relying instead on the user. However, it may require the user to specify a possibly non-trivial test oracle when used beyond input-output examples.
@@ IV-B2 Ranking test suggestions
After obtaining the set of tests produced by an LLM, the user is presented with a sequence of tests. The user response to these proposed tests in both TiCoder scenarios (TiCoder-PassFail, TiCoder-Output) are used to prune and rank code suggestions. To minimize the number of user interactions, it is desirable to prioritize tests that would result in the most number of incorrect code suggestions being pruned away [15, 29]. To achieve this, the set of tests are executed against the set of possible code suggestions generated by the LLM.
Then, using this execution information, we adopt a discriminative test ranking policy that prioritizes tests that can discriminate best among the set of code suggestions generated by the LLM. If a test $t$ can discriminate between code suggestions well (i.e., splits the set of code suggestions into roughly equal halves), then it would prune away a substantial fraction of the code suggestions irrespective of the user response (either Pass or Fail).
More precisely, let $U$ be the set of test suggestions and $G$ be the set of code suggestions that have not been pruned away after $k\geq 0$ user interactions. For each test $t\in U$, we split the set of code suggestions $G$ into the sets $G^{+}_{t}$ and $G^{-}_{t}$ of code suggestions that pass and fail the assertion in $t$, respectively. Note that we ignore codes that results in a crash on a test $t$ instead of failing with an assertion failure. We treat these as precondition violation. We then prioritize tests where the ratio of the sizes of these two set is closest to 1. In other words, we rank the tests in decreasing order using the following scoring metric $s_{discr}$:
@@
Consider the example in Figure 3. Consider the two tests $t_{1}$ and $t_{2}$: Two code suggestions {$c_{2}$, $c_{3}$} FAIL on test suggestion $t_{1}$ while one suggestion {$c_{1}$} PASS, making $s_{\textit{discr}}(t_{1})=\texttt{min}(1,2)/\texttt{max}(1,2)=1/2$. Similarly, two code suggestions {$c_{1}$, $c_{2}$} PASS on test suggestion $t_{2}$ while one suggestion {$c_{3}$} FAIL and $s_{\textit{discr}}(t_{2})=1/2$. All code suggestions in this example PASS on test $t_{3}$ making $s_{\textit{discr}}(t_{3})=0$.
@@ IV-B3 Pruning and ranking code suggestions
TiCoder returns a ranked list of code suggestions, whose behavior is consistent with all the user responses, and prunes the other code suggestions generated by the LLM, whose execution behavior on tests is contradictory to user expectation. Let us first consider the case of code pruning. Let $t\doteq(i,o)$ be a test in the form of an input-output example presented to the user. If the user responds Pass, then we prune any code $c\in G$ for which executing $c(i)\neq o$. Similarly, if the user responds Fail, then we prune any code $c\in G$ for which executing $c(i)=o$. In addition, for TiCoder-Output if the user provides the desired output $o^{\prime}$ for the input $i$, then we can further prune any code suggestion $c$ for which $c(i)\neq o^{\prime}$. Note that we cannot soundly prune any code if the user responds with Undefined.
@@ V-A1 Control condition: AI Programming Assistant 1
Assistant 1 represents the control condition for the experiment. Given the pre-selected prompt, Assistant 1 generates 5 code suggestions for the user, surfaced in a random order. Participants using Assistant 1 always see 5 unique code suggestions. We make this decision to reflect the current user experience scenario of several real-world AI code generation tools, such as GitHub Copilot’s completion panel. For example, the GitHub Copilot completion panel in VSCode shows the user up to 10 possible code suggestions at a time. Our decision is also informed by research pointing to the benefit of surfacing multiple code suggestions[31, 33, 32, 8]. We limit the maximum number of codes to 5 so as to allow the participant to complete each task within 15 minutes.